# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ C.3 Classifying charge assignments for bosonic and spin TQFTs
Specialize first to the bosonic case. Then we can use Thm. C.1 to see that there exists a unique Abelian anyon $v$ such that
@@
In the fermionic/spin case Thm. C.6 means that there exists an Abelian anyon $v$ such that
@@ C.4 From spin TQFT to spinc TQFTs
$\begin{split}Q(1)&=0\text{ mod }{2}\\ Q(c)&=1\text{ mod }{2}\\ Q(x)+Q(y)-Q(z)&=0\text{ mod }{2}\qquad,\qquad\text{ if }N_{xy}^{z}>0\,.\end{split}$ (C.12)
@@
We will present two derivations of the assignment of charges modulo 2 and will explore the properties of the spinc category. Our first approach will be to try generalizing the bosonic approach of App. C.3. We will see that this approach falls short of a full classification, motivating us to employ the bosonic shadow.
@@
We start by following the logic of Thm. C.9, using the consistency of $Q(x)\text{ mod }1$ to expressed it in terms of the vison
@@
Note that here Thm. C.6 means the charge assignment only fixes $v$ up to $c$.
@@
Since these braiding phases should all be ±$\pm 1$, $w^{2}=1$. In other words, the choices are parametrized by a torsor over the 2-torsion subgroup of $\mathcal{A}$. Here torsor means that $\xi$ measures the difference between different charge assignments (all with the same $Q$ mod 1, or the same vison).
@@
We give a somewhat trivial example of assigning charges to a spin TQFT to turn it into a spinc TQFT. Suppose $\cal{T}$ splits, i.e., ${\cal T}={\cal M}\boxtimes\{1,c\}$, where ${\cal M}$ is a bosonic TQFT. Then, one consistent choice is $Q(x)=0$ mod 2 for all anyons in ${\cal M}$, which corresponds to $v=1$. According to our classification, any other charge assignment with the same $Q$ mod 1 is given by
@@
Unlike the bosonic TQFT, in general the vison can not be an arbitrary Abelian anyon due to the spin-charge relation. Below we illustrate the subtlety with the example of ${\cal M}={\rm Ising}$ and provide a complete classification of charge assignments in this theory.
@@
$22t(\sigma)=-2\tilde{B}(v,\sigma)\text{ mod }$ 2 . (C.21)
@@
As the above example reveals, this approach to classifying spinc TQFTs leaves something to be desired. It necessitates manually checking all Abelian anyons to find the choices consistent with being the vison of a spinc TQFT. After this is done, one then needs to find all order two Abelian anyons to finish the classification. This motivates us to consider another approach.
@@
As done in Sec. 2.2, to classify such charge assignment in $\mathcal{T}^{\rm ext}$, we can adiabatically thread $2\pi$-flux ($\pi$-flux) with respect to $Q^{\rm ext}$ ($Q=2Q^{\rm ext}$). As in Sec. 2.2 we identify the result of doing so with an Abelian anyon $\bm{\gamma}$. In other words, $\bm{\gamma}$ is a vison for the bosonic theory ${\cal T}^{\rm ext}$. This anyon must have $\pi$-braiding with $c$ to give $Q^{\rm ext}(c)=\frac{1}{2}\text{ mod }1$, so $\bm{\gamma}\in{\cal T}^{\rm ext}_{1}$. The charge assignment is then completely determined by $\bm{\gamma}$:
$Q(a)=2Q^{\rm ext}(a)$ $=2B(a,\bm{\gamma})\text{ mod }{2}\,.$ (C.22)
@@
$Q(\bm{\gamma})$ $=2Q^{\rm ext}(\bm{\gamma})=2\sigma_{H}^{\rm ext}=\frac{\sigma_{H}}{2}\text{ mod }2$ (C.23)
@@
Let us relate this construction to our first derivation in App. a. The vison $v\in\mathcal{T}^{\rm ext}_{0}=\mathcal{T}$ is determined by
@@
We thus find that $B(v,a)=2B(\bm{\gamma},a)=B(\bm{\gamma}^{2},a)\text{ mod }1$. Thus, we must have $v=\bm{\gamma}^{2}$. Physically, this is obvious; $\bm{\gamma}$ corresponds to $\pi$-flux of $Q$ and $v$ corresponds to $2\pi$-flux of $Q$.
Thus, if the vison in $\cal{T}$ is given, then $\bm{\gamma}^{2}=v$ means that the different choices form a torsor over the subgroup of order-2 elements in $\mathcal{A}$. This is indeed just what we saw in App. a.
@@
Since $\bm{\xi}$ is Abelian, the set $\bm{\xi}\times\mathcal{T}^{\rm ext}_{0}\subseteq{\cal T}^{\rm ext}_{1}$ contains exactly the same number of elements as $\mathcal{T}^{\rm ext}_{0}$. Further, since $\bm{\xi}$ is Abelian $d_{\bm{\xi}_{a}}=d_{a}$. Then the total quantum dimension of $\bm{\xi}\times\mathcal{T}^{\rm ext}_{0}$ is equal to that of ${\cal T}^{\rm ext}_{0}$. Since ${\cal D}_{{\cal T}^{\rm ext}_{0}}={\cal D}_{{\cal T}^{\rm ext}_{1}}$, as proven in App. D, this set saturates the total quantum dimension of $\mathcal{T}^{\rm ext}_{1}$. Thus ${\cal T}^{\rm ext}_{1}=\bm{\xi}\times{\cal T}^{\rm ext}_{0}$.
@@
Note that this fact is consistent with the statement in App. C.2 that the different spin structures are related to each other via continuously tuning boundary conditions. The continuous change in the boundary conditions is the same as the process of flux threading. (In other contexts, it is known as spectral flow.) And the one-to-one map of the anyons is clear by following them during the continuous flux-threading process. In the case of the $\pi$ flux threading, the map (C.25) is with $\bm{\xi}=\bm{\gamma}$, i.e., with the vison of ${\cal T}^{\rm ext}$.
We have shown that charge assignment in the bosonic TQFT $\cal T^{\rm ext}$ such that $Q^{\rm ext}(c)=1/2$ mod 1 is parametrized by ${\cal A}_{1}$. When restricted to $\mathcal{T}$, since anyons in $\mathcal{T}$ braid trivially with $c$, different choices are classified by $\mathcal{A}_{1}/\{1,c\}$. We can show that the classification is complete: that all consistent solutions of $Q(a)$ for $a\in\mathcal{T}$ satisfying the conditions in Definition C.10 are obtained from (C.22) for $\bm{\gamma}\in\mathcal{A}_{1}/\{1,c\}$.
We have now accomplished our goal of classifying the symmetry enrichment of a spinc theory. Thus far, however, our classification relied on a particular choice of a modular extension, ${\cal T}^{\rm ext}$, which possessed an Abelian anyon $\bm{\gamma}\in{\cal T}^{\rm ext}_{1}$. We could accept this, and simply state that the modular extension we chose is the unique one chosen by gauging the fermion parity, as in Rem. C.8. Instead, we show that our classification is independent of this choice. First, we prove two simple corollaries.
@@
Let ${\cal T}$ be a spin TQFT. Then if ${\cal T}^{\rm ext}_{1}$ contains an Abelian anyon, $\bm{\gamma}$, it has no Majorana anyons.
@@
One immediate consequence of these results is that since $\mathcal{T}^{\rm ext}_{1}$ has to contain the Abelian anyon $\bm{\gamma}$, a necessary condition for a spinc TQFT $\mathcal{T}$, is that $\mathcal{T}^{\rm ext}_{1}$ cannot contain any Majorana anyon. This is in accord with the observation made after (C.6), that in a spinc theory $N_{o}=N_{e}$.
@@
Let us first discuss the case of odd $n$ and argue that it is inconsistent with our assumptions. To see that, we will argue that all the $\pi$-flux anyons are non-Abelian. Consider the anyon content of ${\cal T}^{\prime}$ explicitly. We follow the steps outlined in Sec. 3 to gauge the one-form symmetry generated by $\psi c$. First, following steps one and two, the anyons that have trivial braiding with $\psi c$ are either contained in $\mathcal{T}_{0}$ or of the form $\sigma\bm{a}$ for $\sigma\in{\rm Spin}(n)_{1}$, $\bm{a}\in\mathcal{T}^{\rm ext}_{1}$. Next, following step three, we need to determine if any of these anyons are fixed points under fusion with $\psi c$. The anyons in ${\cal T}_{0}$ are clearly not. For the anyons of the form $\sigma\bm{a}$ we have $\sigma\bm{a}\times\psi c=\sigma(\bm{a}\times c)\neq\sigma\bm{a}$ by the results of Corr. C.13 and the fact that ${\cal T}^{\rm ext}$ was chosen such that there is an Abelian anyon in ${\cal T}_{1}^{\rm ext}$. Thus these anyons are not fixed points either and ${\cal T}^{\prime}={\cal T}_{0}^{\prime}\oplus{\cal T}^{\prime}_{1}$ with ${\cal T}_{0}^{\prime}={\cal T}_{0}$ and ${\cal T}_{1}^{\prime}=\{\sigma\bm{\gamma}a\mid\sigma\in{\rm Spin}(n)_{1},a\in\mathcal{T}_{0}/\{1,c\}\}$. Since $\sigma$ is non-Abelian, all of the anyons in ${\cal T}_{1}^{\prime}$ are non-Abelian. Thus, they cannot be the result of $\pi$-flux threading since these anyons must be Abelian, as discussed earlier. Physically, since getting to ${\cal T}^{\prime}$ from ${\cal T}^{\rm ext}$ corresponds to stacking $p+ip$ superconductors, the charge conservation is broken.
Thus, we take $n=2k$. Then, one can write the anyons in Spin$(2k)_{1}$ as $\{1,m,\psi,m\psi\}$ with $B(m,\psi)=1/2,h(m)=k/8$. The anyons in ${\cal T}^{\prime}={\cal T}^{\prime}_{0}\oplus{\cal T}^{\prime}_{1}$ can be described as follows: ${\cal T}^{\prime}_{0}$ is the same as ${\cal T}^{\rm ext}_{0}=\cal{T}$, and ${\cal T}^{\prime}_{1}=m\times{\cal T}_{1}^{\rm ext}$ (thus a one-to-one mapping). Because $m$ is Abelian, there is also a one-to-one mapping between the Abelian subset ${\cal A}_{1}^{\prime}\subset{\cal T}_{1}^{\prime}$, and ${\cal A}_{1}$. In this process $\bm{\gamma}$ is shifted to $m\times\bm{\gamma}$, but this does not affect its braiding with ${\cal T}_{0}={\cal T}_{0}^{\prime}$. Hence, the classification is not affected.
@@
Let us summarize what we have learned. We started with a spinc theory $\cal T$ and derived its minimal bosonic extension ${\cal T}^{\mathrm{ext}}$. [^50] Our discussions up to this point in fact provides an explicit construction of a minimal modular extension $\mathcal{T}^{\rm ext}$, given a spinc charge assignment on $\mathcal{T}$ and the Hall conductivity $\sigma_{H}$. More concretely, $\mathcal{T}^{\rm ext}_{1}$ is given by $\bm{\gamma}\times\mathcal{T}$, where $\bm{\gamma}$ has $h(\bm{\gamma})=\frac{\sigma_{H}}{8}\text{ mod }1$ and $B(a,\bm{\gamma})=\frac{1}{2}Q(a)\text{ mod }1$. Then the full data of the modular extension is determined. This can be done in 8 different ways, labeled by stacking SO$(2k)_{1}$ with $k=0,1,\cdots,7$ before we sum over the spin structures. The difference between these theories is in $\sigma_{H}\to\left(\sigma_{H}+k\right)\text{ mod }8$ and consequently in $h(\bm{\gamma})\to\left(h(\bm{\gamma})+{k\over 8}\right)\text{ mod }1$ and $Q(\bm{\gamma})\to\left(Q(\bm{\gamma})+{k\over 2}\right)\text{ mod }2$. The rest of the modular data is determined up to this $k$ dependence, however, the charge assignment on $\mathcal{T}$ is independent of $k$ [^51] In Sec. 3.3, we argued that in spinc theories, we should identify theories with the same $\sigma_{H}\text{ mod }1$. How come they appear to be different here? The point is that, as we argued in Sec. 3.3, shifting $\sigma_{H}$ by an integer does not change $\cal T$. It affects only contact terms and the behavior of $\cal T$ in the presence of boundaries. Indeed, the Ramond line is an anyon in ${\cal T}^{\mathrm{ext}}$, but it is not an anyon in $\cal T$. One way to think about it in $\cal T$ is by removing a line from our three-manifold and specifying Ramond boundary conditions for the fermions around it, or equivalently, holonomy $-1$ for $A$ around it. This can be thought of as a boundary in $\cal T$.
@@
This quantity was discussed in [167]. It is well-defined for the spinc theory even if $c$ is identified with the vacuum, as reviewed in Sec. 3. It is a quadratic refinement of the topological spin $h$, so the ribbon identity holds.
@@
$\begin{split}e^{\frac{2\pi i}{8}c_{-}}&=\frac{1}{\cal D_{{\cal T}^{\rm ext}}}\left(\sum_{a\in{\cal T}}d_{a}^{2}e^{2\pi ih(a)}+\sum_{\bm{\alpha}\in{\cal T}^{\rm ext}_{1}}d_{\bm{\alpha}}^{2}e^{2\pi ih(\bm{\alpha})}\right)\\ &=\frac{1}{{\cal D}_{{\cal T}^{\rm ext}}}\sum_{\bm{\alpha}\in{\cal T}^{\rm ext}_{1}}d_{\bm{\alpha}}^{2}e^{2\pi ih(\bm{\alpha})}\\ &=\frac{1}{\sqrt{2}{\cal D}_{\cal T}}\sum_{\bm{\alpha}\in{\cal T}^{\rm ext}_{1}}d_{\bm{\alpha}}^{2}e^{2\pi ih(\bm{\alpha})}\\ &=\frac{1}{\sqrt{2}{\cal D}_{\cal T}}\sum_{a\in{\cal T}}d_{a}^{2}e^{2\pi ih(\bm{\gamma}_{a})}\\ &=\frac{1}{\sqrt{2}{\cal D}_{\cal T}}e^{2\pi ih(\bm{\gamma})}\sum_{a\in{\cal T}}d_{a}^{2}e^{2\pi i\hat{h}(a)}\end{split}$ (C.29)
@@
If all anyons carry integer charges, then by Thm. C.6 the vison $v$ is either $1$ or $c$.
If $v=1$, then we must have $\bm{\gamma}^{2}=1$. We also have $B(\bm{\gamma},c)=1/2$. Thus we find that the theory $\{1,\bm{\gamma},c,\bm{\gamma}c\}$ is already an MTC. [^52] There are four options given by Spin$(4k)_{1}$ for $k=0,1,2,3$. By Thm C.2, ${\cal T}^{\rm ext}$ must factorize as ${\cal T}^{\prime}\boxtimes\{1,\bm{\gamma},c,\bm{\gamma}c\}$ where $\cal T^{\prime}$ is an MTC. Then we see that ${\cal T}={\cal T}^{\rm ext}_{0}={\cal T}^{\prime}\boxtimes\{1,c\}$.
We can repeat this argument for $\bm{\gamma}^{2}=c$, in which case the subcategory $\{1,\bm{\gamma},c,\bm{\gamma}c\}$, is also modular. [^53] There are four options given by Spin$(4k+2)_{1}=\mathcal{A}^{4,2k+1}$ for $k=0,1,2,3$.
@@
Consider a split spinc theory, $\mathcal{T}={\cal M}\boxtimes\{1,c\}$, with an MTC $\cal M$, where all the anyons have integer charge. Then, there exists an MTC ${\cal M}^{\prime}$ such that $\mathcal{T}={\cal M}^{\prime}\boxtimes\{1,c\}$, and the anyons in ${\cal M}^{\prime}$ have zero charges mod 2.
@@
We will directly construct ${\cal M}^{\prime}$. The anyon types are in one-to-one correspondence with the anyon types of ${\cal M}$, through the following map:
@@
It is clear that $\{x^{\prime}\}$ follow the same fusion rules of $\cal M$, i.e., $N^{x^{\prime}y^{\prime}}_{z^{\prime}}=N^{xy}_{z}$.
@@
It is straightforward to check that the pentagon and hexagon identities are still satisfied. The modular data of ${\cal M}^{\prime}$ is given by
@@
We finish by using Thm. C.7 to write ${\cal T}={\cal M}^{\prime}\boxtimes{\cal T}^{\prime}$, where ${\cal T}^{\prime}$ is a super-modular category. Since ${\cal D}_{{\cal M}^{\prime}}={\cal D}_{{\cal M}}={\cal D}_{{\cal T}}/\sqrt{2}$, this mean ${\cal D}_{{\cal T}^{\prime}}=\sqrt{2}$. Then, we conclude ${\cal T}^{\prime}=\{1,c\}$. ∎
Corollary C.19.
If $\sigma_{H}\in\mathbb{Z}$, the super-modular category must split. Equivalently, a non-split spinc theory has a fractional $\sigma_{H}$.
Proof.
If $\sigma_{H}$ is even, then the vison $v$ is an even charged boson, while if it is odd, then $vc$ is. Then one can gauge the one-form symmetry generated by this even charged boson, while preserving the $\mathrm{U}(1)$ symmetry. The resulting theory is still spinc, but now with a trivial vison. Then by Thm. C.17 it takes the form of ${\cal M}^{\prime}\boxtimes\{1,c\}$. One can use the previous corollary to define ${\cal M^{\prime}}$ in such a way that the dual zero-form symmetry acts just on ${\cal M}^{\prime}$. Then the original super-modular category can be recovered by gauging the dual zero-form symmetry, which proves that it also splits. ∎
@@
Note that the same result was found in App. C.2.

### v3

@@ C.3 Classifying charge assignments for bosonic and spin TQFTs
Specialize first to the bosonic case. Then we can use Theorem C.1 to see that there exists a unique Abelian anyon $v$ such that
@@
In the fermionic/spin case Theorem C.6 means that there exists an Abelian anyon $v$ such that
@@ C.4 From spin TQFT to spinc TQFTs
[…]
@@
We will present two derivations of the assignment of charges modulo 2 and will explore the properties of the spinc category. Our first approach will be to try generalizing the bosonic approach of Appendix C.3. We will see that this approach falls short of a full classification, motivating us to employ the bosonic shadow.
@@
We start by following the logic of Theorem C.9, using the consistency of $Q(x)\text{ mod }1$ to expressed it in terms of the vison
@@
Note that here Theorem C.6 means the charge assignment only fixes $v$ up to $c$.
@@
Since these braiding phases should all be $\pm 1$, $w^{2}=1$. In other words, the choices are parametrized by a torsor over the 2-torsion subgroup of $\mathcal{A}$. Here torsor means that $\xi$ measures the difference between different charge assignments (all with the same $Q$ mod 1, or the same vison).
@@
[…]
@@
Unlike the bosonic TQFT, in general the vison can not be an arbitrary Abelian anyon owing to the spin-charge relation. Below we illustrate the subtlety with the example of ${\cal M}={\rm Ising}$ and provide a complete classification of charge assignments in this theory.
@@
$2t(\sigma)=-2\tilde{B}(v,\sigma)\text{ mod }2.$ (C.21)
@@
As the above example reveals, this approach to classifying spinc TQFTs leaves something to be desired. It necessitates manually checking all Abelian anyons to find the choices consistent with being the vison of a spinc TQFT. After this is done, one needs to find all order two Abelian anyons to finish the classification. This motivates us to consider another approach.
@@
As done in Sec. 2.2, to classify such charge assignment in $\mathcal{T}^{\rm ext}$, we can adiabatically thread $2\pi$ flux ($\pi$ flux) with respect to $Q^{\rm ext}$ ($Q=2Q^{\rm ext}$). As in Sec. 2.2 we identify the result of doing so with an Abelian anyon $\bm{\gamma}$. In other words, $\bm{\gamma}$ is a vison for the bosonic theory ${\cal T}^{\rm ext}$. This anyon must have $\pi$ braiding with $c$ to give $Q^{\rm ext}(c)=\frac{1}{2}\text{ mod }1$, so $\bm{\gamma}\in{\cal T}^{\rm ext}_{1}$. The charge assignment is then completely determined by $\bm{\gamma}$,
[…]
@@
[…]
@@
Let us relate this construction to our first derivation in Appendix a. The vison $v\in\mathcal{T}^{\rm ext}_{0}=\mathcal{T}$ is determined by
@@
We thus find that $B(v,a)=2B(\bm{\gamma},a)=B(\bm{\gamma}^{2},a)\text{ mod }1$. Thus, we must have $v=\bm{\gamma}^{2}$. Physically, this is obvious; $\bm{\gamma}$ corresponds to $\pi$ flux of $Q$ and $v$ corresponds to $2\pi$ flux of $Q$.
Thus, if the vison in $\cal{T}$ is given, then $\bm{\gamma}^{2}=v$ means that the different choices form a torsor over the subgroup of order-2 elements in $\mathcal{A}$. This is indeed just what we saw in Appendix a.
@@
Since $\bm{\xi}$ is Abelian, the set $\bm{\xi}\times\mathcal{T}^{\rm ext}_{0}\subseteq{\cal T}^{\rm ext}_{1}$ contains exactly the same number of elements as $\mathcal{T}^{\rm ext}_{0}$. Further, since $\bm{\xi}$ is Abelian $d_{\bm{\xi}_{a}}=d_{a}$. Then the total quantum dimension of $\bm{\xi}\times\mathcal{T}^{\rm ext}_{0}$ is equal to that of ${\cal T}^{\rm ext}_{0}$. Since ${\cal D}_{{\cal T}^{\rm ext}_{0}}={\cal D}_{{\cal T}^{\rm ext}_{1}}$, as proven in Appendix D, this set saturates the total quantum dimension of $\mathcal{T}^{\rm ext}_{1}$. Thus ${\cal T}^{\rm ext}_{1}=\bm{\xi}\times{\cal T}^{\rm ext}_{0}$.
@@
Note that this fact is consistent with the statement in Appendix C.2 that the different spin structures are related to each other via continuously tuning boundary conditions. The continuous change in the boundary conditions is the same as the process of flux threading. (In other contexts, it is known as spectral flow.) And the one-to-one map of the anyons is clear by following them during the continuous flux-threading process. In the case of the $\pi$ flux threading, the map (C.25) is with $\bm{\xi}=\bm{\gamma}$, i.e., with the vison of ${\cal T}^{\rm ext}$.
[…]
We have now accomplished our goal of classifying the symmetry enrichment of a spinc theory. Thus far, however, our classification relied on a particular choice of a modular extension ${\cal T}^{\rm ext}$, which possessed an Abelian anyon $\bm{\gamma}\in{\cal T}^{\rm ext}_{1}$. We could accept this, and simply state that the modular extension we chose is the unique one chosen by gauging the fermion parity, as in Remark C.8. Instead, we show that our classification is independent of this choice. First, we prove two simple corollaries.
@@
Let $\cal T$} be a spin TQFT. Then if ${\cal T}^{\rm ext}_{1}$ contains an Abelian anyon $\bm{\gamma}$, it has no Majorana anyons.
@@
One immediate consequence of these results is that since $\mathcal{T}^{\rm ext}_{1}$ has to contain the Abelian anyon $\bm{\gamma}$, a necessary condition for a spinc TQFT $\mathcal{T}$ is that $\mathcal{T}^{\rm ext}_{1}$ cannot contain any Majorana anyon. This is in accord with the observation made after (C.6), that in a spinc theory $N_{o}=N_{e}$.
@@
Let us first discuss the case of odd $n$ and argue that it is inconsistent with our assumptions. To see that, we will argue that all the $\pi$-flux anyons are non-Abelian. Consider the anyon content of $\cal T$$^{\prime}$ explicitly. We follow the steps outlined in Sec. 3 to gauge the one-form symmetry generated by $\psi c$. First, following steps one and two, the anyons that have trivial braiding with $\psi c$ are either contained in $\mathcal{T}_{0}$ or of the form $\sigma\bm{a}$ for $\sigma\in{\rm Spin}(n)_{1}$, $\bm{a}\in\mathcal{T}^{\rm ext}_{1}$. Next, following step three, we need to determine if any of these anyons are fixed points under fusion with $\psi c$. The anyons in ${\cal T}_{0}$ are clearly not. For the anyons of the form $\sigma\bm{a}$ we have $\sigma\bm{a}\times\psi c=\sigma(\bm{a}\times c)\neq\sigma\bm{a}$ by the results of Corrollary C.13 and the fact that ${\cal T}^{\rm ext}$ was chosen such that there is an Abelian anyon in ${\cal T}_{1}^{\rm ext}$. Thus these anyons are not fixed points either and ${\cal T}^{\prime}={\cal T}_{0}^{\prime}\oplus{\cal T}^{\prime}_{1}$ with ${\cal T}_{0}^{\prime}={\cal T}_{0}$ and ${\cal T}_{1}^{\prime}=\{\sigma\bm{\gamma}a\mid\sigma\in{\rm Spin}(n)_{1},a\in\mathcal{T}_{0}/\{1,c\}\}$. Since $\sigma$ is non-Abelian, all of the anyons in ${\cal T}_{1}^{\prime}$ are non-Abelian. Thus, they cannot be the result of $\pi$-flux threading since these anyons must be Abelian, as discussed earlier. Physically, since getting to $\cal T$$^{\prime}$ from ${\cal T}^{\rm ext}$ corresponds to stacking $p+ip$ superconductors, the charge conservation is broken.
[…]
@@
[…]
@@
This quantity was discussed in [165]. It is well defined for the spinc theory even if $c$ is identified with the vacuum, as reviewed in Sec. 3. It is a quadratic refinement of the topological spin $h$, so the ribbon identity holds.
@@
[…]
@@
[…]
Now by Theorem C.2, ${\cal T}^{\rm ext}$ must factorize as ${\cal T}^{\prime}\boxtimes\{1,\bm{\gamma},c,\bm{\gamma}c\}$ where $\cal T^{\prime}$ is an MTC. Recall that ${\cal T}^{\rm ext}={\cal T}\oplus{\cal T}\times\bm{\gamma}$, where the elements in $\cal T$} are precisely those that braid trivially with $c$. It is then not hard to see that since ${\cal T}^{\rm ext}={\cal T}^{\prime}\boxtimes\{1,\bm{\gamma},c,\bm{\gamma}c\}$ for some MTC $\cal T$$^{\prime}$, then the elements braiding trivially with $c$ are precisely ${\cal T}^{\prime}\times\{1,c\}$. So then ${\cal T}={\cal T}^{\prime}\times\{1,c\}$ and we see our super-modular category must split. ∎
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
We finish by using Theorem C.7 to write ${\cal T}={\cal M}^{\prime}\boxtimes{\cal T}^{\prime}$, where $\cal T$$^{\prime}$ is a super-modular category. Since ${\cal D}_{{\cal M}^{\prime}}={\cal D}_{{\cal M}}={\cal D}_{{\cal T}}/\sqrt{2}$, this mean ${\cal D}_{{\cal T}^{\prime}}=\sqrt{2}$. Then, we conclude ${\cal T}^{\prime}=\{1,c\}$. ∎
@@
Note that the same result was found in Appendix C.2.